OpenStudy (anonymous):

if \[f(x)=kx^3-1 \] and \[f ^{-1}(15)=2\], find k.

OpenStudy (whpalmer4):

What does \(f^{-1}(15) = ?\)

OpenStudy (anonymous):

= 2 sorry!

OpenStudy (whpalmer4):

I would figure out the formula for \(f^{-1}(x)\) in the usual fashion of swapping x and y and then solving for y. Then plug in the known value and solve for k.

OpenStudy (anonymous):

\[f^{-1}(15)=\frac{ 1 }{ kx^3-1 }\] \[2=\frac{ 1 }{ k(15)^3-1 }\]

OpenStudy (whpalmer4):

@Azteck that's supposed to be the inverse function, I think, not the reciprocal of f(x).

OpenStudy (anonymous):

Oh crap. yeah sorry about that. It's meant to be this. \[f^{-1}(15)=\sqrt[3]{\frac{ y-1 }{ k }}\]

OpenStudy (anonymous):

x-1 sorry*

OpenStudy (anonymous):

not y-1

OpenStudy (whpalmer4):

Too bad it isn't \(f(x) = 1/x\) as \(f^{-1}(x) = 1/f(x)\) in that case :-)

OpenStudy (anonymous):

oh okay so uhmm... i got \[f^{-1}(x) = \sqrt[3]{\frac{ x+1 }{ 3 }}\]

OpenStudy (anonymous):

Yeah that's right @alexeis_nicole I made a mistake on not changing the sign.

OpenStudy (anonymous):

sorry... i meant k not 3 as a denominator

OpenStudy (anonymous):

\[2=\sqrt[3]{\frac{ 15+1 }{ k }}\] \[8=\frac{ 15+1 }{ k }\] \[8k=16\] \[k=2\]

OpenStudy (anonymous):

@whpalmer4 @Azteck THANK YOU !

OpenStudy (anonymous):

You're welcome.

OpenStudy (whpalmer4):

I tell you how to do it, and point out when the other guy is leading you astray, but he gets the medal. Harrumph :-)

OpenStudy (whpalmer4):

Even if @Azteck hadn't given me a medal, I'd still be a happy camper, as you two gave me the idea of finding the function whose inverse is its reciprocal, don't think I ever noticed that before!